OpenStudy (iwik):

Help !!?!?!?!?!?! Given the following functions f(x) and g(x), solve f[g(6)] and select the correct answer below: f(x) = 6x + 12 g(x) = x − 8 A) −96 B)0 C)24 D)48

OpenStudy (anonymous):

You need to first substitute 6 for x in the g(x) equation What would that get you?

OpenStudy (iwik):

g(x)=2

rebeccaxhawaii (rebeccaxhawaii):

no it would be -2

rebeccaxhawaii (rebeccaxhawaii):

because 6-8=-2

rebeccaxhawaii (rebeccaxhawaii):

then plug -2 into the f equation

OpenStudy (iwik):

ok so f(x) = 6-2 + 12 =f(x)=16

rebeccaxhawaii (rebeccaxhawaii):

incorrect

OpenStudy (iwik):

wait f(x)=0?

rebeccaxhawaii (rebeccaxhawaii):

6(-2)+12 -12+12=0 therefore f(g(x))=0 because you plug x into the g(x) equation first then your g(x)= -2 so then you plug -2 into the f(x) then that gives you your final answer

rebeccaxhawaii (rebeccaxhawaii):

Please click "best response" for the person that helped you the most, this just shows the appreciation for the time they took to assist. If you want further help with anymore questions, please fan me so in the long run you can see when I'm on. I'll be sure to take a look at your question to help :)

OpenStudy (iwik):

ok thank 4 ur help :)